OpenStudy (anonymous):

Can someone help me solve this question using the guess method? A popular radio station broadcasts with a frequency of 105.1 MHz. What is the wavelength of the broadcast? (1 M10Hz=10^6Hz)

OpenStudy (phi):

Guessing the answer sounds hard (impossible?) I would use freq * wavelength = speed (of light) 105.1*10^6 per sec* L = 3*10^8 m/sec you get \[ L = \frac{3\cdot 10^8}{105.1\cdot 10^6} m \] which can be simplified.

OpenStudy (phi):

\[ 105.1 \cdot 10^6 = 1.051 \cdot 10^8\] (if you make the 105 smaller by dividing by 100, you make the 10^6 bigger by multiplying by 100) so the wavelength is \[ L = \frac{3\cdot 10^8}{1.051\cdot 10^8} \ meters\] or very close to 3 meters (about 10 feet)

OpenStudy (wolf1728):

wavelength = speed of light / frequency wavelength = 299,792,458 meters per second / 105,100,000

OpenStudy (anonymous):

Thanks - that makes it a little more clear!! :)